OpenStudy (anonymous):

factor completely and show steps will get medal 50k3 – 40k² + 75k – 60

OpenStudy (mertsj):

Factor out 5 first.

OpenStudy (anonymous):

ok

OpenStudy (promisespromises):

50k3 - 40k2 + 75k - 60 5(10K3-8k2+15k-12) 5[(10k3-8k2) + (15K-12)] 5[2k2 (5k-4) + 3(5k-4)] 5[(2k2+3)+(5k-4)] I believe that's the final answer.

OpenStudy (anonymous):

wow thx

OpenStudy (promisespromises):

Welcome :)

OpenStudy (anonymous):

anyone think they can help with another one

OpenStudy (promisespromises):

I'll give it a shot if you want

OpenStudy (anonymous):

find the volume and surface area of the cube 4x-5

OpenStudy (anonymous):

5(2k^2+3)(5k-4)

OpenStudy (promisespromises):

Let 4x-5 = a Volume = a^3 V= (4x-5)^3 And Surface area = 6a^2 SA= 6(4x-5)^2

OpenStudy (anonymous):

thx a million

OpenStudy (promisespromises):

No problemo